### v1

@@ Abstract
The experimental results are included in the following video: https://youtu.be/75AbKY3L5As.
@@ 1 Introduction
(a) Training procedure of MFRL method.
(b) The proposed planning policy and reward estimator.
Figure 2: (a) Overview of the proposed training method. The method employs multi-fidelity deep reinforcement learning to train a policy focused on minimizing trajectory time duration. Rewards for reinforcement learning are estimated by a reward estimator, which efficiently models the feasibility boundary using multi-fidelity policy evaluations. (b) Overview of the planning policy and reward estimator architecture. The planning policy, based on a sequence-to-sequence model, determines the flight time between remaining waypoints. The reward estimator predicts feasibility of the policy outputs using a multi-fidelity Gaussian process classifier.
@@
This paper addresses the challenge of time-optimal online trajectory planning for agile vehicles, a vital component in various unmanned aerial vehicle (UAV) applications. Crafting high-speed quadrotor trajectories is particularly demanding due to the complex aerodynamics, turning trajectory generation into a time-consuming non-linear optimization. Moreover, real-world constraints, including the effects of control delay and state estimation error, can make the generated trajectory unreliable even if it satisfies the ideal dynamics constraints. Online planning adds yet another layer of complexity by incorporating computational time constraints to update the trajectory in real-time. As a consequence, many existing online planning methods rely on simplified constraints such as bounding velocity and acceleration Gao et al. (2018a); Tordesillas et al. (2019); Wu et al. (2023), or on hierarchical planning structures that combine the approximated global planner with the underlying control schemes Romero et al. (2022a), which may lead to conservative solutions.
@@
The main contribution of this paper is a novel multi-fidelity reinforcement learning (MFRL) framework. MFRL integrates RL and BO to develop a planning policy optimized for real-world and real-time scenarios. BO efficiently models the feasibility boundary of time-optimal trajectories by selectively evaluating uncertain policy outputs through precise high-fidelity evaluations, including complex simulations and real-world tests. We further improve the modeling efficiency by combining the evaluations with low-fidelity evaluations—a lightweight approach using a simple dynamics model—via the multi-fidelity technique (multi-fidelity Bayesian optimization; MFBO). Concurrently, RL optimizes the policy to minimize the duration of output trajectories within these feasibility boundaries. The trained policy is validated through comprehensive numerical simulations and real-world flight experiments. It outperforms the baseline minimum-snap method by producing trajectories that are not only faster and more reliable but also generated with significantly lower computational time. Specifically, the output trajectories achieve up to a 25 % reduction in flight time, averaging a 4.7 % decrease, with just 2 ms of computational time—markedly less than the several minutes required by the baseline.
This framework is unique for the following reasons: Firstly, we directly model the feasibility boundary of the time-optimal trajectories using BO. In prior work, the complex dynamics model could not be easily integrated with the RL training procedure, and they relied on simplified kinodynamics constraints like velocity and acceleration limits, resulting in suboptimal performance. Our method significantly enhances planning policy performance by accurately modeling the feasibility boundary, where time-optimal trajectories are typically found. Secondly, our use of MFBO streamlines the modeling cost of trajectory feasibility boundaries, traditionally requiring extensive high-fidelity experiments. By incorporating low-fidelity analytical assessments through the MFBO framework, we significantly reduce the number of high-fidelity experiments required. Thirdly, the policy trained via MFRL enables real-time generation of time-optimal trajectories for scenarios demanding online planning. The BO-derived feasibility boundary model, due to its computational demands for uncertainty prediction, is not suited for real-time inference. We instead utilize it for reward estimation during RL training, ensuring the resulting RL policy is computationally efficient and optimized for real-time applications.
The proposed method is built upon the MFBO framework used in Ryou et al. (2021, 2022). In Ryou et al. (2021), MFBO is employed to create trajectories for specific waypoint sequences, while in Ryou et al. (2022), they expand this to general waypoint sequences by pretraining with a MFBO-labeled subdataset and fine-tuning through reinforcement learning. This current work introduces significant advancements, particularly in training efficiency and real-time re-planning. We integrate MFBO with the RL training process, effectively shortening training time and enabling real-world experiment inclusion, contrasting the previous simulation-only approach. The model proposed in this study uses the current trajectory state as an input to the planning policy, enabling trajectory planning from non-stationary states, crucial for re-planning in mid-course. Additionally, our planning model directly determines the absolute time for waypoint traversal, unlike the previous method’s use of proportional time distributions. This eliminates the need for extra line-search procedures to maintain trajectory feasibility, thereby enhancing the computational efficiency of trajectory generation.
@@ 2.1 Quadrotor Trajectory Planning
Quadrotor planning often employs snap-minimization, leveraging the differential flatness of quadrotor dynamics for smooth trajectory generation Mellinger and Kumar (2011); Richter et al. (2016). Minimizing the fourth-order derivative of position, i.e. , snap, and the yaw acceleration produces a smooth trajectory that is less likely to violate feasibility constraints. Several methods that outperform the snap-minimization method have been presented in recent years. Foehn et al. (2021) employs a general time-discretized trajectory representation, using waypoint proximity constraints to find better, time-optimal solutions. Gao et al. (2018b) streamlines minimum-snap computation by partitioning the optimization problem, separating speed profile and trajectory shape optimizations. Sun et al. (2021); Burke et al. (2020) employ spatial optimization gradients to solve non-linear optimization for time allocation, which offers numerical stability advantages over naive non-linear optimization. Qin et al. (2023) improves the time optimization by incorporating linear collision avoidance constraints into the objective function through projection methods, transforming the problem into unconstrained optimization. Wu et al. (2023) utilizes supervised learning to learn time allocation, minimizing the computation time for time optimization. These approaches optimize trajectory time while adhering to the snap minimization objective. Mao et al. (2023), however, directly minimizes time using reachability analysis within speed and acceleration bounds. Instead of using the predefined velocity and acceleration constraints, Ryou et al. (2021) utilizes MFBO to model the feasibility constraints from the data and utilizes the model to directly minimize trajectory time.
@@
Most offline trajectory generation methods, due to the limited computation time, require approximation to be extended to the online replanning problem. Gao et al. (2018a) formulates the online planning problem as quadratic programming that can be solved in real-time by directly optimizing the velocity and acceleration that conforms to the specified velocity and acceleration constraints. Tordesillas et al. (2019) uses the time allocations from the previous trajectory to warm-start the optimization for the updated trajectory, thereby reducing the computation time. Romero et al. (2022a) builds a velocity search graph between the prescribed waypoints sequence and finds the optimal velocity profile with Dijkstra search. As the trajectory is generated in discrete state space, the resulting trajectory may not be feasible, so feasible control inputs are generated using underlying model predictive control Romero et al. (2022b) or deep neural network policy Molchanov et al. (2019). On the other hand, Kaufmann et al. (2023) trains a neural network policy to directly generate thrust and body rate commands based on the vehicle’s position using reinforcement learning, achieving superior performance even surpassing human racers. Nonetheless, further research is needed to expand the applicability of this learning-based approach to deploy the trained policy effectively in unseen, diverse environments.
@@ 2.2 Efficient Reinforcement Learning
Uncertainty-aware RL is an additional strategy to boost sample efficiency, which uses uncertainty estimation to streamline exploration throughout the RL training phase. Osband et al. (2016) uses bootstrapping method to estimate policy output variance from dataset subsets, enhancing sample efficiency by selecting outputs that maximize rewards from the policy outputs’ posterior distribution. Kumar et al. (2019) uses the same method of uncertainty estimation to evaluate out-of-distribution samples, aiming to stabilize the training process. Wu et al. (2021) utilizes MC-dropout, which is introduced in Gal and Ghahramani (2016), to predict the variance of predictions and regularizes the training error with the inverse of this variance. Similarly, Clements et al. (2019); Lee et al. (2021) apply an ensemble method, which involves using multiple policy models to estimate the uncertainty, facilitating safer exploration during reinforcement learning. Reviews of these methodologies are included in Lockwood and Si (2022); Hao et al. (2023). To handle high-fidelity samples with very limited data, we adopt a Gaussian process framework with a multi-fidelity kernel, moving away from neural-network-based uncertainty estimation. We instead compensate for the limited model capacity of the Gaussian process using a variational approach, elaborated on in the subsequent section.
@@
Several methods have been proposed to deploy trained models in real-world robotic systems, addressing the discrepancies between simulated and real-world environments. One approach involves introducing random perturbations to the input of policy (Sadeghi et al. (2018); Akkaya et al. (2019)) or dynamics model (Peng et al. (2018); Mordatch et al. (2015)) during RL training. This method produces robust policies capable of handling errors related to the simulation-to-reality gap. However, this may compromise the policy’s performance, making it unsuitable for tasks that demand precise modeling of the system’s limits, like time-optimal trajectory generation. Alternatively, the real-world deployment can be enhanced by increasing the accuracy of the simulation. For instance, using system identification techniques, studies such as Tan et al. (2018); Kaspar et al. (2020); Hwangbo et al. (2019) incorporate real-world constraints like actuation delay and friction into simulations. Meanwhile, Chang and Padif (2020); Lim et al. (2022); Chebotar et al. (2019) iteratively refine simulations based on real-world data using learning-based methods. Nevertheless, these approaches often overlook the costs of real-world experiments, which involve significant and costly evaluations to address the disparity between simulation and reality. Furthermore, as simulations become more accurate, their evaluation costs rise proportionally, rendering them less suitable for RL training that demands millions of training samples.
@@ 3.1 Minimum-Snap Trajectory Planning
The snap-minimization method utilizes the smooth polynomial that is obtained by minimizing the fourth-order derivative of position and the yaw acceleration. For a sequence of waypoints $\tilde{p}^{i}$, each consisting of a prescribed position $\tilde{p}^{i}_{r}\in\mathbb{R}^{3}$ and yaw angle $\tilde{p}^{i}_{\psi}\in\mathbb{S}^{1}$ ($\tilde{\mathbf{p}}=\begin{bmatrix}\tilde{p}^{0}\cdots\tilde{p}^{m}\end{bmatrix}$), this approach models the trajectory using piecewise continuous polynomials. These polynomials map time to position and yaw, i.e. , $p(t)=\begin{bmatrix}{{}p_{r}(t)}^{\intercal}&p_{\psi}(t)\end{bmatrix}^{\intercal}$ ($p_{r}(t)\in\mathcal{C}^{4},p_{\psi}(t)\in\mathcal{C}^{2}$), and each polynomial segment connects two adjacent waypoints. $\mathbb{S}^{1}$ denotes the circular angle space, and $\mathcal{C}^{n}$ is the differentiability class where its n-th order derivatives exist and are continous.
@@ 3.2 Multi-Fidelity Gaussian Process
In this paper, we employ the Multi-fidelity Gaussian Process Classifier (MFGPC) to efficiently model the feasibility boundary based on sparse multi-fidelity evaluations. Given a collection of data points $\mathbf{Z}=\{\mathbf{z}_{1},\cdots,\mathbf{z}_{N}\}$ with their associated evaluation outcomes from the $l$-th fidelity level $\mathbf{y}_{l}=\{y_{l,1},\cdots,y_{l,N}\}$, the prior Gaussian distribution is modeled using a multi-fidelity covariance kernel. The key idea of the multi-fidelity covariance kernel $K_{l}(\mathbf{Z},\mathbf{Z})$ is utilizing cheap low-fidelity evaluations to efficiently model the high-fidelity measurements. The relationship between the distributions of different fidelity levels is represented by the nonlinear space-dependent transformation proposed in Cutajar et al. (2019). To be specific, an element of $l$-th covariance kernel $k_{l}(z,z^{\prime})$ is estimated as
@@
where $\mathcal{B}(x)$ is the Bernoulli likelihood and $\Phi(f_{l,i})$ is the cumulative density function used to map the latent variable $f_{l,i}$ onto the probability domain [0, 1]. The prediction on the l-th fidelity level over the new data points is obtained by marginalizing the latent variable as follows:
[…]
@@
[…]
@@
The losses incurred at lower fidelity levels are added to provide regularization, with $c_{\text{reg}}$ representing the coefficient for the regularization term. During the MFRL training, samples with high uncertainty estimates from the boundary model are selected, enabling iterative multi-fidelity evaluations and boundary model updates. MFGPC is implemented with GPyTorch Gardner et al. (2018) based on the work presented in Ryou et al. (2021).

### v2

@@ Abstract
The experimental results are included in the following video: https://youtu.be/75AbKY3L5As. Detailed implementation and source code are available at: https://github.com/mit-aera/mfrlTrajectory.
@@ 1 Introduction
This paper addresses the challenge of time-optimal online trajectory planning for agile vehicles with dynamic waypoint changes, as illustrated in Figure 1. This capability represents a vital component in various unmanned aerial vehicle (UAV) applications. However, crafting high-speed quadrotor trajectories is particularly demanding due to the complex aerodynamics, turning trajectory generation into a time-consuming non-linear optimization. Moreover, real-world constraints, including the effects of control delay and state estimation error, can make the generated trajectory unreliable even if it satisfies the ideal dynamics constraints. Online planning adds yet another layer of complexity by incorporating computational time constraints to update the trajectory in real-time. As a consequence, many existing online planning methods rely on simplified constraints such as bounding velocity and acceleration Gao et al. (2018a); Tordesillas et al. (2019); Wu et al. (2023), or on hierarchical planning structures that combine the approximated global planner with the underlying control schemes Romero et al. (2022a), which may lead to conservative solutions.
@@
The main contribution of this paper is a novel multi-fidelity reinforcement learning (MFRL) framework. MFRL integrates RL and BO to develop a planning policy optimized for real-world and real-time scenarios. First, this method utilizes BO to directly model feasibility boundaries from high-fidelity evaluations—accurate but expensive methods, e.g., real-world experiments. It improves policy performance by accurately modeling constraint boundaries where time-optimal trajectories are typically found. Second, we implement multi-fidelity Bayesian optimization (MFBO) to reduce required high-fidelity evaluations by using low-fidelity evaluations—quick but less accurate methods, e.g., analytical checks with ideal dynamics—to rapidly screen out excessively fast or slow trajectories with obvious tracking results. Thirdly, RL uses this constraint modeling in the reward signal for policy training, resulting in a computationally efficient policy that enables real-time generation of time-optimal trajectories for online planning scenarios. Lastly, the trained policy is validated through comprehensive numerical simulations and real-world flight experiments. It outperforms the baseline minimum-snap method by achieving up to a 25 % reduction in flight time, averaging a 4.7 % decrease, with just 2 ms of computational time—markedly less than the several minutes required by the baseline.
@@ 2.1 Quadrotor Trajectory Planning
Quadrotor planning often employs snap minimization, leveraging the differential flatness of quadrotor dynamics for smooth trajectory generation Mellinger and Kumar (2011); Richter et al. (2016). Minimizing the fourth-order derivative of position, i.e. , snap, and the yaw acceleration produces a smooth trajectory that is less likely to violate feasibility constraints. Several methods that outperform the snap minimization method have been presented in recent years. Foehn et al. (2021) employs a general time-discretized trajectory representation, using waypoint proximity constraints to find better, time-optimal solutions. Gao et al. (2018b) streamlines minimum-snap computation by partitioning the optimization problem, separating speed profile and trajectory shape optimizations. Sun et al. (2021); Burke et al. (2020) employ spatial optimization gradients to solve non-linear optimization for time allocation, which offers numerical stability advantages over naive non-linear optimization. Qin et al. (2023) improves the time optimization by incorporating linear collision avoidance constraints into the objective function through projection methods, transforming the problem into unconstrained optimization. Wu et al. (2023) utilizes supervised learning to learn time allocation, minimizing the computation time for time optimization. These approaches optimize trajectory time while adhering to the snap minimization objective. Mao et al. (2023), however, directly minimizes time using reachability analysis within speed and acceleration bounds. Instead of using the predefined velocity and acceleration constraints, Ryou et al. (2021) utilizes MFBO to model the feasibility constraints from the data and utilizes the model to directly minimize trajectory time.
@@
Most offline trajectory generation methods, due to the limited computation time, require approximation to be extended to the online replanning problem. Gao et al. (2018a) formulates the online planning problem as quadratic programming that can be solved in real-time by directly optimizing the velocity and acceleration that conforms to the specified velocity and acceleration constraints. Tordesillas et al. (2019) uses the time allocations from the previous trajectory to warm-start the optimization for the updated trajectory, thereby reducing the computation time. Romero et al. (2022a) builds a velocity search graph between the prescribed waypoints sequence and finds the optimal velocity profile with Dijkstra search. As the trajectory is generated in discrete state space, the resulting trajectory may not be feasible, so feasible control inputs are generated using underlying model predictive control Romero et al. (2022b) or deep neural network policy Molchanov et al. (2019). On the other hand, Kaufmann et al. (2023) trains a neural network policy to directly generate thrust and body rate commands based on the vehicle’s position using RL, achieving superior performance even surpassing human racers. Nonetheless, further research is needed to expand the applicability of this learning-based approach to deploy the trained policy effectively in unseen, diverse environments.
These online planning methods often employ local planning with discrete state space optimization, allowing for greater expressiveness and easier implementation of constraints. In contrast, our paper opts for a polynomial trajectory representation. We choose this representation for its reduced optimization dimension, facilitating lighter neural networks and easier dataset creation for machine learning. Additionally, this approach allows for longer planning horizons, incorporating more future information for better overall performance. While the polynomial representation has trade-offs compared to discrete approaches, it has potential to improve expressiveness through additional variables like knot points or higher-degree polynomials.
@@ 2.2 Efficient Reinforcement Learning
Uncertainty-aware RL is an additional strategy to boost sample efficiency, which uses uncertainty estimation to streamline exploration throughout the RL training phase. Osband et al. (2016) uses bootstrapping method to estimate policy output variance from dataset subsets, enhancing sample efficiency by selecting outputs that maximize rewards from the policy outputs’ posterior distribution. Kumar et al. (2019) uses the same method of uncertainty estimation to evaluate out-of-distribution samples, aiming to stabilize the training process. Wu et al. (2021) utilizes MC-dropout, which is introduced in Gal and Ghahramani (2016), to predict the variance of predictions and regularizes the training error with the inverse of this variance. Similarly, Clements et al. (2019); Lee et al. (2021) apply an ensemble method, which involves using multiple policy models to estimate the uncertainty, facilitating safer exploration during RL. Reviews of these methodologies are included in Lockwood and Si (2022); Hao et al. (2023). To handle high-fidelity samples with very limited data, we adopt a Gaussian process framework with a multi-fidelity kernel, moving away from neural-network-based uncertainty estimation. We instead compensate for the limited model capacity of the Gaussian process using a variational approach, elaborated on in the subsequent section.
@@
Several methods have been proposed to deploy trained models in real-world robotic systems, addressing the discrepancies between simulated and real-world environments. One approach involves introducing random perturbations to the input of policy (Sadeghi et al. (2018); Akkaya et al. (2019)) or dynamics model (Peng et al. (2018); Mordatch et al. (2015)) during RL training. This method produces robust policies capable of handling errors related to the simulation-to-reality gap. However, this may compromise the policy’s performance, making it unsuitable for tasks that demand precise modeling of the system’s limits, like time-optimal trajectory generation. Alternatively, the real-world deployment can be enhanced by increasing the accuracy of the simulation. For instance, using system identification techniques, studies such as Tan et al. (2018); Kaspar et al. (2020); Hwangbo et al. (2019) incorporate real-world constraints like actuation delay and friction into simulations. Meanwhile, Chang and Padif (2020); Lim et al. (2022); Chebotar et al. (2019) iteratively refine simulations based on real-world data using learning-based methods. These approaches improve simulation accuracy and computational efficiency, making them more suitable for RL training. However, some components, like battery dynamics, remain challenging to model accurately. Moreover, complex systems such as autonomous vehicles often can’t be comprehensively simulated. These limitations necessitate real-world experiments on robot deployment, which our paper aims to make more efficient.
@@ 3.1 Minimum-Snap Trajectory Planning
The snap minimization method utilizes the smooth polynomial that is obtained by minimizing the fourth-order derivative of position and the second-order derivative of yaw. For a sequence of waypoints $\tilde{p}^{i}$, each consisting of a prescribed position $\tilde{p}^{i}_{r}\in\mathbb{R}^{3}$ and yaw angle $\tilde{p}^{i}_{\psi}\in\mathbb{S}^{1}$ ($\tilde{\mathbf{p}}=\begin{bmatrix}\tilde{p}^{0}\cdots\tilde{p}^{m}\end{bmatrix}$), this approach models the trajectory using piecewise continuous polynomials. These polynomials map time to position and yaw, i.e. , $p(t)=\begin{bmatrix}{{}p_{r}(t)}^{\intercal}&p_{\psi}(t)\end{bmatrix}^{\intercal}$ ($p_{r}(t)\in\mathcal{C}^{4},p_{\psi}(t)\in\mathcal{C}^{2}$), and each polynomial segment connects two adjacent waypoints. $\mathbb{S}^{1}$ denotes the circular angle space, and $\mathcal{C}^{n}$ is the differentiability class where its n-th order derivatives exist and are continous.
@@ 3.2 Multi-Fidelity Gaussian Process
In this paper, we employ the Multi-fidelity Gaussian Process Classifier (MFGPC) to efficiently model the feasibility boundary based on sparse multi-fidelity evaluations. Given a collection of data points $\mathbf{Z}=\{\mathbf{z}_{1},\cdots,\mathbf{z}_{N}\}$ with their associated evaluation outcomes from the $l$-th fidelity level $\mathbf{y}_{l}=\{y_{l,1},\cdots,y_{l,N}\}$, the prior Gaussian distribution is modeled using a multi-fidelity covariance kernel. The key idea of the multi-fidelity covariance kernel $K_{l}(\mathbf{Z},\mathbf{Z})$ is utilizing cheap low-fidelity evaluations to efficiently model the high-fidelity measurements. The relationship between the distributions of different fidelity levels is represented by the nonlinear space-dependent transformation proposed in Cutajar et al. (2019). To be specific, $k_{l}(z,z^{\prime})$, an element of the $l$-th covariance kernel $K_{l}(\mathbf{Z},\mathbf{Z})$ corresponding to data points $z$ and $z^{\prime}$, is estimated as
@@
Figure 2: Multi-fidelity Gaussian process classifier training and inferencing. Training: Lower fidelity kernel captures the overall structure using ample data to optimize more hyperparameters. Higher fidelity learns linear transformation on top. Covariance between dataset pairs, $z$ and $z^{\prime}$, (Cyan line) is estimated using previous level guidance. Inferencing: Covariances estimated between existing and a new data $z^{*}$ (Red line) are used for prediction via marginalization of training variables.
@@
where $\mathcal{B}(x)$ is the Bernoulli likelihood and $\Phi(f_{l,i})$ is the cumulative density function used to map the latent variable $f_{l,i}$ onto the probability domain [0, 1]. The prediction, $y_{*}$, on the l-th fidelity level over a new data point, $z_{*}$, is obtained by marginalizing the latent variable, $f_{*}$, as follows:
$P(y_{*}|\mathbf{z}_{*},\mathbf{Z},\mathbf{y})=\int P(y_{*}|f_{*})P(f_{*}|\mathbf{z}_{*},\mathbf{Z},\mathbf{y}_{l})d\mathbf{f_{*}}.$ (9)
Figure 2 illustrates the training and inferencing procedure of the multi-fidelity Gaussian process classifier.
@@
where $q(\mathbf{f})\coloneqq\int P(\mathbf{f}|\mathbf{u})q(\mathbf{u})d\mathbf{u}$. $\mathcal{M}\coloneqq\left(\mathbf{f},\theta,\mathbf{m},\mathbf{S}\right)$ is the set of all hyperparameters, and $\theta$ is the hyperparameters of the Gaussian process kernel. The hyperparameters of the MFGPC are determined by minimizing the sum of the variational lower bounds across all fidelity levels:
@@
The losses incurred at lower fidelity levels, $\mathcal{L}_{\mathcal{M}_{l^{\prime}}}$ where $l^{\prime}\in\{1,\dots,l-1\}$, are added to provide regularization, with $c_{\text{reg}}$ representing the coefficient for the regularization term. During the MFRL training, samples with high uncertainty estimates from the boundary model are selected, enabling iterative multi-fidelity evaluations and boundary model updates. MFGPC is implemented with GPyTorch Gardner et al. (2018) based on the work presented in Ryou et al. (2021).
3.3 Comparison with Prior Work
The proposed method is built upon the MFBO framework used in Ryou et al. (2021, 2022). In Ryou et al. (2021), MFBO is employed to create trajectories for specific waypoint sequences, while in Ryou et al. (2022), they expand this to general waypoint sequences by pretraining with a MFBO-labeled subdataset and fine-tuning through RL. This current work introduces significant advancements, particularly in training efficiency and real-time re-planning. We integrate MFBO with the RL training process, effectively shortening training time and enabling real-world experiment inclusion, contrasting the previous simulation-only approach. The model proposed in this study uses the current trajectory state as an input to the planning policy, enabling trajectory planning from non-stationary states, crucial for re-planning in mid-course. Additionally, our planning model directly determines the absolute time for waypoint traversal, unlike the previous method’s use of proportional time distributions. This eliminates the need for extra line-search procedures to maintain trajectory feasibility, thereby enhancing the computational efficiency of trajectory generation.